Math 135 - section 01 - Precalculus Diagnostic Quiz August 30, 2013 Your Name: Show your work and circle the correct answer. Please turn over for problems 4,5,6.

1. Let f(x) = x^{2} − x and g(x) = x + 1. Which function is f(g(x))? A. x^{2} − x B. x^{2} + 3x C. x^{2} + x D. x^{2} + x + 2

2. Find common factors and cancel to simplify: \frac{12x}{3x − 6} · \frac{x^{2} − 4}{2x + 4}

A. \frac{x^{2} + 12x − 4}{5x − 2} B. 2x C. −\frac{4}{3}(x − 1) D. \frac{12x^{3} − 48x}{6x^{2} − 12}

3. Which is equal to (u^{−5}v^{2})^{3} (\frac{v^{2}}{u} )^{−1}?

A. u^{−16}v^{8} B. u^{−14}v^{4} C. u^{−9}v^{3} D. (uv)^{−7} 4. If f(x) = 5x^{2} − 11, what is f(a + 1) − f(a)?

5. Find all real number values of x satisfying 2(x − 2) > 5.

6. Solve for x: 2x^{2} − x − 6 = 0 (find all real number solutions).